A space station, in the form of a wheel 118 m in diameter, rotates to provide an "artificial gravity" of 2.90 m/s2 for persons who walk around on the inner wall of the outer rim. Find the rate of the wheel's rotation in revolutions per minute that will produce this effect.
? rev/min
Answer #1

radius, r= 118/2= 59 m

expression for acceleration in rotating motion

a= r*w^2

2.9= 59*w^2

w= 0.2217 rad/sec

w= 2.118 rev/min

plz rate